Super bikes arrive in a 12 lot showroom. Inspections are carried out on six out of every 12. For
one lot, it is known 2 out of 12 do not meet the safety standards prescribed. What is the
probability that at least 2 out of the 6 that were tested from that lot will not meet safety
standards?
1
Expert's answer
2020-06-15T19:05:42-0400

Poisson distribution

P(x=X)=(λxe)/x!

P(x≥2)=P(x=2)+P(x=3)+P(x=4)+P(x=5)+P(x=6)=(22e-2)/2!+(23e-2)/3!+(24e-2)/4!+(25e-2)/5!+(26e-2)/6!=0.27+0.18+0.09+0.036+0.012=0.588